OpenStudy (anonymous):

Can anyone please tell me how to differentiate between a chain rule and a power rule ? I don't know which rule to apply when I see a mixed equation

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

the power rule is for a single function; the chain rule is for a function inside of a function

OpenStudy (amistre64):

f(x) = x^2 ; power rule: 2x f(g(x)) = sqrt(x^2+1); chain rule: 2x/2sqrt(x^2+1)

OpenStudy (amistre64):

whenever you an split the function into nested function; more than one thing happens; you apply the chain rule

OpenStudy (anonymous):

yes but sometimes its not easy to see...for example ....(1-7t) to the power of 6, the prof told us to apply chain rule and power rule

OpenStudy (amistre64):

the chain rule is what you apply; but the power rule is part of the setup

OpenStudy (amistre64):

(1-7t)^6 is a function in a function: (....)^6 and 1-7t; derive both

OpenStudy (amistre64):

6(1-7t)^5 (-7) = -42(1-7t)^5

OpenStudy (anonymous):

ok how about sin ( tan square root sinx)

OpenStudy (amistre64):

sin(tan(sqrt(sin(x)))) that?

OpenStudy (amistre64):

sin(...); tan(....); sqrt(....); sin(.....); x derive it thru now

OpenStudy (anonymous):

ok so this is a chain rule I know but the tan and square root of sin x is that a product rule ?

OpenStudy (amistre64):

sin(...); tan(....); sqrt(....); sin(.....); x cos(tan(sqrt(sin(x)))); sec^2(sqrt(sin(x))); 1/2sqrt(sin(x)); cos(x); 1

OpenStudy (amistre64):

if you split it out onto like a conveyor belt; you can just derive its parts

OpenStudy (amistre64):

and the chain rule says when your done; multiply them all together

OpenStudy (amistre64):

\[\frac{cos(tan(\sqrt{sin(x)}))*sec^2(\sqrt{sin(x)})*cos(x)}{2\sqrt{sin(x)}}\]

OpenStudy (anonymous):

ok I just want to know what rule u applied inside ( tan square root of sinx)

OpenStudy (anonymous):

I understand how u differentiate but I just want to know the rule inside the ( )

OpenStudy (amistre64):

the chain rule to split it up; and the only 'power' rule would have been to take care of the sqrt

OpenStudy (amistre64):

you have to peel it like an onion; then work on it from the inside out

OpenStudy (amistre64):

you cant just pick and choose which ones you do and dont; its all one unit that needs to be accounted for

OpenStudy (amistre64):

\[sin(tan(\sqrt{sin(x)}))\] \[sin(...)*tan(....)*\sqrt{(....)}*sin(....)*x\] \[cos(...)*sec^2(...)*\frac{1}{2\sqrt{(....)}}*cos(...)*1\]

OpenStudy (amistre64):

then you place it all back together like a set of russian dolls

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

thanx a lot

OpenStudy (amistre64):

:) youre welcome

OpenStudy (anonymous):

I wish u were my prof :) Thanx again